Assume that Whee, Cheatham, and Howe is an auditing firm that has found that its summer interns are subject to a 50 percent lTime required for 4th financial statement Time required for thinancial statement Time required for 16th financial statementPercentage of proofreading cost for 16th financial statement to cost of proofreading ist financial statement
Answer #1

Solution:

(a) The standard equation for the learning curve: Tn=T1nb

Here,

n is the number of financial proofread

Tn = Amount of time required to read n proofreads

T1 = Amount of time required to read first financial proofread

b = learning curve factor which is equals to ln(p) / ln(2)

here ln(p) is the natural logarithm pf p

p =percentage of learning (in this case it is 50%)

Using the information given in the question:

b = ln(0.5) / ln(2) = -1

Number of financial proofreads T1nb The time required to read hours * 60

1

6*(1-1) 6 hours 360 minutes
2 6*(2-1) 3 hours 180 minutes
4 6*(4-1) 1.5 hours 90 minutes
8 6*(8-1) 0.75 hours 45 minutes
16 6*(16-1) 0.375 hours 22.5 minutes

(b) The labor time cost is $11 per hour.

Cost of financial proofreading = Cost per hour * Number of hours required to read

Cost of proofreading 1st financial statement = $11 * 6 hours = $66

Cost of proofreading 16th financial statement = $11 * 0.375 hours = $4.125

The percentage of proofreading cost for 16th year to cost of proofreading for 1st year =

= Cost of proofreading 16th financial statement / Cost of proofreading 1st financial statement

= $4.125 / $66

= 0.0625 = 6.25%